OpenStudy (anonymous):

a collection of 30 coins worth $5.50 consist of nickels, dimes and quarters, rhere are twice as many dimes as nickels ,how many quarters are there

OpenStudy (amistre64):

n + d + q = 30 get all the values in terms of one variable; the variable they give you is "n". n = n d = 2n q = 30-n-2n = 30-3n remember these relationships; they come in handy at the end. We know this: n(.05) + 2n(.10) + (30-3n)(.25) = 5.50 ; solve for "n".

OpenStudy (amistre64):

.05n +.20n -.75n + 7.50 = 5.50 combine like terms. -.50n = -2.00 ; multiply each side by (-1) .50n = 2.00 ; divide each side by (.50) n = 2/.5 = 20/5 = 4 n=4 d=2(4) = 8 q=30 - 4 - 8 = 30-12 = 18